OpenStudy (nkhanfar):

can someone help me determine if y=-x^2-1 makes y=0

OpenStudy (mrs.ambrose614):

I awnsered your other quetsion for you now hold on while I figure this one out

OpenStudy (will.h):

could you demonstrate what it means idk what they mean is there an x value you need to plug in and find if the result would be y = 0?

OpenStudy (nkhanfar):

it asks detirmine if the parbola opens up or down (up a larger than 0 down a less than 0 a is -1 and the axis of symmytry

OpenStudy (mrs.ambrose614):

Use the slope form y=Mx+b to find the y intercept b b=0

OpenStudy (mrs.ambrose614):

Y=-x^2-1 makes y=0 I hope I helped you

OpenStudy (nkhanfar):

so axis is on (0, 0)?

OpenStudy (mrs.ambrose614):

Yes it is

OpenStudy (nkhanfar):

Thanks so much!

OpenStudy (mrs.ambrose614):

I used an app called math way to find the awnser

OpenStudy (will.h):

alright y=-x^2-1 this is an even degree parabola the leading coefficent is negative thus the parabola would be upside down the axis of symmetry can be found using -b/2a the equation needs to be adjusted by the order of degrees thus y = -x^2 + 0x - 1 -b/2a -0/-2 axis of symmetry would be zero hope that helps